OpenStudy (hoolol2):

Subtract. g-2/6 - g+3/9

OpenStudy (hoolol2):

I think the answer is g-12/18

OpenStudy (hoolol2):

am I correct?

OpenStudy (fwizbang):

not quite. Add the g's and the numbers separately.

OpenStudy (hoolol2):

is g-5/18 correct?

OpenStudy (hoolol2):

@fwizbang

OpenStudy (fwizbang):

still not quite there. Just to be clear, does the problem ask you to subtract g+3/9 from g-2/6 or is it (g+3)/9 from (g-2)/6?

OpenStudy (hoolol2):

there are no parenthesis in the problem

OpenStudy (fwizbang):

Ok, so we start with g-2/6 and subtract g-3/9. Treat the g's separately from the numbers. So what is g-g?

OpenStudy (hoolol2):

0

OpenStudy (fwizbang):

good. Now -2/6 =?

OpenStudy (hoolol2):

-1/3?

OpenStudy (fwizbang):

Good! 3/9=?

OpenStudy (hoolol2):

1/3

OpenStudy (fwizbang):

Terrific.So (-2/6)-3/9 = (-1/3) - (1/3) = ?

OpenStudy (hoolol2):

I think so

OpenStudy (fwizbang):

So what's the final answer?

OpenStudy (hoolol2):

-2/3?

OpenStudy (fwizbang):

Perfect!

OpenStudy (hoolol2):

so is -2/3 the final answer

OpenStudy (hoolol2):

@fwizbang

OpenStudy (fwizbang):

yes, well done.

OpenStudy (hoolol2):

ok thanks